You have been hired to replace the manager of We-Fix-Any-Car, a firm that uses only two inputs, capital and labor, to produce output. The firm can hire as much labor, as it wants at a wage of $500 per day, per worker and can rent as much capital as it wants at a price of $25,000 per day. After you look at the company books, you learn that the company has been using capital and labor in amounts that imply a marginal product of the labor of 6,000 and a marginal product of capital of 100,000. Do you know why the firm, We-Fix-Any-Car let the last manager go and hired you? Explain.
Answer #1

The optimal condition for the input is given by:

MPL/MPK = w/r

6000/100000 = 500/25000

0.06 is not equal to 0.02

Since the inputs are not being employed optimally so We fix any car let the manager go and employed me as the Manager to employ the inputs to minimize the costs